# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

### v1

@@ Abstract
We propose a new asymptotic equipartition property for the perplexity of a large piece of text generated by a language model and present theoretical arguments for this property. Perplexity, defined as a inverse likelihood function, is widely used as a performance metric for training language models. Our main result states that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that language models are constrained to only produce outputs from a “typical set”, which we show, is a vanishingly small subset of all possible grammatically correct outputs. We present preliminary experimental results from an open-source language model to support our theoretical claims. This work has possible practical applications for understanding and improving “AI detection” tools and theoretical implications for the uniqueness, predictability and creative potential of generative models.
@@ 1.1 Contribution: Equipartition Property for Perplexity
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister and Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover and Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg and Keller, 2008; Cohen and Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory. The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara and Morita, 2000; Timo et al., 2010). The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
@@ 2.1 Sequential LLM
$P({\mathbf{Y}}_{N}|{\mathbf{X}})$ $=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
@@
The perplexity $p_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(p_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
@@ 2.2 A Toy Problem
where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the relative frequency of token $y$ in the (long) string ${\mathbf{A}}_{N}$. The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
@@
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(\alpha)~{}{\mathrm{with~{}equality~{}iff~{}}}p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
Proposition 1.
@@
Proposition 1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e. convergence in probability. applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
$l_{M}({\mathbf{A}}_{N})$ $\rightarrow E\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
@@
$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(\alpha)\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961).
2.3 A Naive Method for Detecting Text Generated by a Language Model
Given a text string ${\mathbf{Y}}$, we seek to determine whether or not it was generated by a given model M. This is the problem of “white-box” detection of AI generated text where we assume knowledge of the specific language model that potentially generated the text. A naive method for doing this is as follows: compare the log-perplexity $l_{M}({\mathbf{Y}})$ of the text ${\mathbf{Y}}$ for model M with a threshold: if $l_{M}({\mathbf{Y}})<l_{th}$, we determine that ${\mathbf{Y}}$ is likely to have been generated by model M.
Consider a long string ${\mathbf{B}}\doteq[B_{1},~{}B_{2},~{}\dots,~{}B_{N}]$ randomly generated from model B. Using the same argument as in Section 2.2, we have:
$l_{B}({\mathbf{B}})$ $\equiv H(p_{\mathbf{B}},\beta)\rightarrow H(\beta)\leq H(p_{\mathbf{B}},\alpha)\equiv l_{A}({\mathbf{B}})$ (10)
Note, however, that we have no basis to assert any relationship between $l_{A}({\mathbf{A}})$ and $l_{A}({\mathbf{B}})$. Indeed if $H(\beta)\ll H(\alpha)$ i.e. the distribution $\beta$ has significantly lower entropy than $\alpha$, it cannot be ruled out that $l_{A}({\mathbf{B}})<l_{A}({\mathbf{A}})$ i.e. a text-string generated by a different model has lower perplexity for a model than a text-string generated by itself.
The mathematical properties of cross-entropy provide a rigorous basis for using perplexity to compare the compatibility of different models with a given text string. However, this theory does not support using perplexity to compare different text strings for compatibility with a given model.
@@ 2.4 A Modest Generalization
This theoretical analysis relies on Proposition 1 which only applies to toy models that generate strings of iid tokens. Of course, the simple AEP in Proposition 1 has been extended in various ways in the literature such as the following version (see Tal et al. (2017), Appendix B.4), which we will make use of in the sequel.
Proposition 2.
@@
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,~{}\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (11)
@@
While Proposition 2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable. Much of the analysis in Section 2.2 can be extended to generalized models like A’ in Proposition 2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$. However, the tokens in these models must be drawn from fixed distributions independent of past tokens. This is still quite trivial compared to modern language models where the output tokens depend in highly complex and sophisticated ways on past tokens.
@@ 3 An Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.2 to a more interesting class of models. Consider a language model M and a random infinitely long text string ${\mathbf{Y}}$ generated by M, whose probabilities for a given prompt ${\mathbf{X}}$ is described by (2). Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$. We will assume that the prompt ${\mathbf{X}}$ is fixed.
Definition 1.
@@
$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{M}H(p_{n})$ (12)
@@
Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number. Indeed, we can show that
$E\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv E\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$ (13)
@@
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}}~{}\mathrm{where}~{}S(p)\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (14)
@@
We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes.
Preview of our main result. We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. But this is very wrong - while the log-deviation $\lambda(p)$ of a fixed probability distribution $p(y)$ is a deterministic number, $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ is a random variable that takes different numerical values for different random text strings ${\mathbf{Y}}$. This considerably complicates our analysis. Fundamentally, their randomness arises from the fact that the distributions $p_{n}(y)$ are different for different strings ${\mathbf{Y}}$. This is because the tokens $Y_{n}$ in the string ${\mathbf{Y}}$ are not generated independently; instead future tokens strongly depend on earlier tokens. This is an essential property of modern language models that allows them to generate sophisticated text outputs. We now introduce the conceptual device of an auxiliary language model to get around this difficulty.
3.1 Concept of an Auxiliary Language Model
Let us consider a fixed string ${\mathbf{Y}}$. For this particular string ${\mathbf{Y}}$ and model M, we have a fixed set of distributions $p_{n}(y)$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{Y}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $p_{n}(y)$. This construct of the auxiliary model $M^{\prime}_{\mathbf{Y}}$ for a fixed string ${\mathbf{Y}}$ allows us to work with fixed distributions $p_{n}(y)$ which are not themselves functions of a random string.
In particular, note that the empirical entropy $h_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ and log-deviation $\lambda_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ for model $M^{\prime}_{\mathbf{Y}}$ are both deterministic functions of $N$ independent of the string ${\mathbf{Y}}^{\prime}_{N}$. Specifically, we have $h_{M^{\prime}_{\mathbf{Y}}}(N)\equiv h_{M}({\mathbf{Y}}_{N})$. Similarly, we have $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)\equiv\lambda_{M}({\mathbf{Y}}_{N})$ which, since we have fixed our special string ${\mathbf{Y}}$, is now also fixed. Indeed, it is easy to show that $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)$ is the standard-deviation of the random variable $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ the log-likelihood of the random string ${\mathbf{Y}}^{\prime}_{N}$. Then, by the Chebyshev Inequality (Cohen, 2015), we have:
$\Pr\left(\left|l_{M}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{\mathbf{Y}}}(N)\right|>\epsilon\right)$ $\leq\frac{\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)}{\epsilon^{2}}$ (16)
Clearly, for substrings ${\mathbf{Y}}_{N}$ of our special string ${\mathbf{Y}}$ that defines the auxiliary model $M^{\prime}_{\mathbf{Y}}$, we have by definition $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N})$. In addition, the entropy $H(P({\mathbf{Y}}^{\prime}_{N}))$ of the random strings ${\mathbf{Y}}^{\prime}_{N}$ output by auxiliary model $M^{\prime}_{\mathbf{Y}}$ is by definition $H(P({\mathbf{Y}}^{\prime}_{N}))\equiv h_{M}({\mathbf{Y}}_{N})$.
The auxiliary model $M^{\prime}_{\mathbf{Y}}$ is much simpler than the general language model M. In particular, because the tokens $Y^{\prime}_{n}$ are generated independently of other tokens, the AEP in Proposition 2 applies to strings ${\mathbf{Y}}^{\prime}$ generated by the auxiliary model $M^{\prime}_{\mathbf{Y}}$. Indeed we note from (15) that the log-deviation $\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)$ vanishes for large $N$ under some regularity conditions on the distributions $p_{n}(y)$, in which case, the AEP for model $M^{\prime}_{\mathbf{Y}}$ follows from (16).
3.2 Ensemble of Auxiliary Models
We can think of a generative model M as a statistical ensemble of a large number of (simpler) auxiliary models. Each auxiliary model is capable of generating a random text string consisting of a sequence of independent tokens. We can then imagine choosing randomly from one of the auxiliary models. We now make this idea more precise.
Consider an enumeration ${\mathbf{Y}}^{(1)},~{}{\mathbf{Y}}^{(2)},~{}\dots,~{}\dots,~{}{\mathbf{Y}}^{(N|\mathcal{T}|)}$ of the possible text strings ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$. Let $M_{{\mathbf{Y}}^{(n)}}$ denote the auxiliary model associated with model M and string ${\mathbf{Y}}^{(n)}$. A sequential generative model can be represented by a probability tree where each branch in the tree represents a new token being generated. The model M as well as each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can each be represented by such a probability tree. An example of a probability tree representing a generative model and a sample random string ${\mathbf{Y}}$ represented by a sequence of branches through the tree are shown in Fig. 1 which also shows the probability tree for one possible auxiliary model $M^{\prime}_{\mathbf{Y}}$.
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens. Also shown in red is the probability tree of an auxiliary model for a string ${\mathbf{Y}}$ with $Y_{1}=3,~{}Y_{2}=1$ and so on.
A generative model uses a pseudo-random number generator (PRNG) that produces a sequence of random symbols ${\mathbf{W}}$ to randomly choose a sequence of branches through its probability tree. As an example, suppose ${\mathbf{W}}=[W_{1},~{}W_{2},~{}\dots W_{N}~{}\dots]$ where the $W_{n}$ are independent and uniformly distributed in $[0,1]$ and a generative model uses $W_{n}$ to generate the n’th token $Y_{n}$ according to the distribution $p_{n}(y)$ using a sampling algorithm such as the classical inversion method Martino et al. (2018). Once we fix the PRNG and the sampling method, we can in principle write down a deterministic function to describe the generative model M as ${\mathbf{Y}}_{N}=m({\mathbf{X}},{\mathbf{W}}_{N})$ where recall that ${\mathbf{X}}$ is the fixed initial prompt. Each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can also similarly be described by deterministic functions ${\mathbf{Y}}_{N}=m_{{\mathbf{Y}}^{(n)}}({\mathbf{X}},{\mathbf{W}}_{N})$.
We can then define a new generative model that we call the ensemble model $M_{ensemble}$ as follows. Pick a pseudo-random sequence ${\mathbf{W}}_{1}$ and generate a string ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$ from model M. Now consider the auxiliary model $M_{{\mathbf{Y}}_{1}}$ corresponding to string ${\mathbf{Y}}_{1}$. Generate a pseudo-random string ${\mathbf{Y}}_{2}=m_{{\mathbf{Y}}_{1}}({\mathbf{X}},{\mathbf{W}}_{2})$ from model $M_{{\mathbf{Y}}_{1}}$ where ${\mathbf{W}}_{2}$ is another pseudo-random sequence. The ensemble model $M_{ensemble}$ is defined by ${\mathbf{Y}}_{2}=m_{ensemble}({\mathbf{X}},{\mathbf{W}}_{1},{\mathbf{W}}_{2})$.
The following observation is crucial to our main result: a good random number generator will almost always generate pseudo-random sequences ${\mathbf{W}}$ that belong to a typical set i.e. ${\mathbf{W}}$ that satisfies the simple AEP in Proposition 1.
Definition 4.
Typical String. An infinite text string ${\mathbf{Y}}$ (or more precisely a sequence of text strings ${\mathbf{Y}}_{N},~{}N=1,2,\dots$) is said to be typical for a model M if the log-perplexity of sufficiently long sub-strings converges to their empirical entropy for model M i.e. if $\lim_{N\rightarrow\infty}|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|=0$.
We illustrate these ideas for a simple Markov language model in Section A.2.
Proposition 3.

### v3

@@ Abstract
We prove a new asymptotic equipartition property for the perplexity of long texts generated by a language model and present supporting experimental evidence from open-source models. Specifically we show that the logarithmic perplexity of any large text generated by a language model must asymptotically converge to the average entropy of its token distributions. This defines a “typical set” that all long synthetic texts generated by a language model must belong to. We show that this typical set is a vanishingly small subset of all possible grammatically correct outputs. These results suggest possible applications to important practical problems such as (a) detecting synthetic AI-generated text, and (b) testing whether a text was used to train a language model. We make no simplifying assumptions (such as stationarity) about the statistics of language model outputs, and therefore our results are directly applicable to practical real-world models without any approximations.
Language Models, Perplexity, Asymptotic Equipartition, Typical Sets
@@ 1.1 Contribution: Equipartition Property for Perplexity
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister & Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover & Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg & Keller, 2008; Cohen & Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory. The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara & Morita, 2000; Timo et al., 2010). The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
Our main contribution is to formulate and provide theoretical and experimental justification for a version of the AEP that applies to real-world practical LLMs without making simplifying assumptions about their statistics. A secondary contribution is to show possible applications of this theory to practical problems such as AI text detection and LLM dataset inference.
@@ 2.1 Sequential LLM
$P$ $({\mathbf{Y}}_{N}|{\mathbf{X}})=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),$
$~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
@@
The perplexity $\mathrm{perp}_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(\mathrm{perp}_{M}({\mathbf{Y}}_{N})\right)$
$\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
@@ 2.2 A Toy Problem
where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the number of occurrences of token $y$ in the (long) string ${\mathbf{A}}_{N}$. The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))$
$\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
@@
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(p_{\mathbf{A}})$
${\mathrm{with~{}equality~{}iff~{}}}$ $p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover & Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
Proposition 2.1.
@@
Proposition 2.1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e. convergence in probability. applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
$l_{M}({\mathbf{A}}_{N})$ $\rightarrow\operatorname*{{E}}\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
@@
$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(p_{\mathbf{A}})$
$\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided only that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961). This same reasoning also underlies popular tools for LLM-related applications as we discuss in Section 5.
@@ 2.3 A Modest Generalization
This theoretical analysis relies on Proposition 2.1 which only applies to toy models that generate strings of iid tokens. Of course, the simple AEP in Proposition 2.1 has been extended in various ways in the literature such as the following version (see (Tal et al., 2017), Appendix B.4), which we will make use of in the sequel.
Proposition 2.2.
@@
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,$
$\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (10)
@@
While Proposition 2.2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable. Much of the analysis in Section 2.2 can be extended to generalized models like $A^{\prime}$ in Proposition 2.2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$. However, the tokens in these models must be drawn from fixed distributions independent of past tokens. This is still quite trivial compared to modern language models where the output tokens depend in highly complex and sophisticated ways on past tokens.
@@ 3 An Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.2 to a more interesting class of models. Consider a language model M and a random infinitely long text string ${\mathbf{Y}}$ generated by M, whose probabilities for a given prompt ${\mathbf{X}}$ is described by (2). Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$. We will assume that the prompt ${\mathbf{X}}$ is fixed and omit it from our notation in the sequel.
Definition 3.1.
@@
$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{N}H(p_{n})$ (11)
@@
Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number. Indeed, we can show that
$\operatorname*{{E}}\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv\operatorname*{{E}}\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N})\right)$ (12)
@@
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}},$
$\mathrm{where}~{}S(p)$ $\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (13)
@@
We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes. We are now ready to state our main result.
Proposition 3.1.